OpenStudy (anonymous):

a = 7 in, b=11 in, c = 5 in. Use Heron's formula to find the area of the triangle.

OpenStudy (anonymous):

T = \sqrt{s(s-a)(s-b)(s-c)} s=(7+11+5)/2 s = 11.5 sbstituting \sqrt{11.5(11.5-7)(11.5-11)(11.5-5) =\sqrt 168.1875 =12.9687.............